# Adaptive Incentive Design with Learning Agents

### v2

@@ 3.1 Fixed point analysis
$\langle\nabla\tilde{\Phi}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger})),\tilde{x}-{\tilde{x}^{\ast}}(\tilde{p}^{\dagger})\rangle\geqslant 0\quad\forall\ \tilde{x}\in\tilde{X}.$ (16)
Recall from the first order conditions of optimality for social cost function, we know that $\tilde{x}^{\dagger}$ is socially optimal if any only if it satisfies
$\langle\nabla\tilde{\Phi}(\tilde{x}^{\dagger}),\tilde{x}-\tilde{x}^{\dagger}\rangle\geqslant 0,\quad\forall\ \tilde{x}\in\tilde{X}.$ (17)
Comparing (16) with (17), we note that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is the minimizer of social cost function $\tilde{\Phi}$. This implies ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})={\tilde{x}}^{\dagger}$ as ${\tilde{x}}^{\dagger}$ is the unique minimizer of social cost function $\tilde{\Phi}$ by Assumption 2.2.
Next, we show that the set $P^{\dagger}$ is singleton. We prove this via contradiction. Suppose $P^{\dagger}$ contains two element $p^{\dagger}_{1},p^{\dagger}_{2}$, which are shown to align the Nash equilibrium with social optimality. Therefore, it holds that $x^{\dagger}=x^{\ast}(p^{\dagger}_{1})=x^{\ast}(p^{\dagger}_{2})$. Moreover, due to (LABEL:eq:_PDagger) we know that $p^{\dagger}_{1}=e(x^{\ast}(p^{\dagger}_{1}))$ and $p^{\dagger}_{2}=e(x^{\ast}(p^{\dagger}_{2}))$. Consequently, we have $p^{\dagger}_{1}=e(x^{\dagger})=p^{\dagger}_{2}$. The proof of uniqueness of $\tilde{P}^{\dagger}$ also follows analogously. ∎
Remark 3.2.
Proposition 3.1 demonstrates that the externality-based incentive updates ($p$-update) and ($\tilde{p}$-update) have two advantages. Firstly, these updates guarantee that any fixed point must achieve social optimality. Secondly, the result indicates that the optimal incentive mechanism takes a simple format: the incentive mechanism is linear in players’ strategies in atomic games and is a constant vector in non-atomic games.
@@ 3.2 Convergence to optimal incentive mechanism
In this subsection, we provide sufficient conditions for convergence of strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update). In order to study the convergence of these updates we make the following assumption:
@@
The step sizes in ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) satisfy the following condition
@@
    $\lim_{k\to\infty}\frac{\beta_{k}}{\gamma_{k}}=0$.
In Assumption 3.1, (i) is a standard assumption on step-sizes that allows us to analyze the convergence of the discrete-time learning updates via a continuous-time differential equation [Borkar, 1997]. Additionally, (ii) assumes that the incentive update occurs at a slower timescale compared to the update of strategies.
In view of Assumption 3.1-(ii), it is necessary that the strategy updates converge if the incentives are held fixed. This condition holds true for a variety of updates commonly studied in learning in games literature. For instance, the equilibrium update (cf. (4)) converges to Nash equilibrium in one-step. Similarly, the best response update (cf. (5)) converges in zero sum games [Harris, 1998], potential games [Swenson et al., 2018], and dominance solvable games [Nachbar, 1990]. The gradient-based update (cf. (6)) converge in concave games [Rosen, 1965]. Therefore, in this work we do not focus on studying the convergence of strategy updates with fixed incentive mechanisms and we pick any off-the-shelf convergent strategy update. Our goal is to characterize conditions under which the coupled strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) converge. Therefore, we impose the following assumption which is sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms. Particularly, we assume that if the incentive mechanism is fixed, then the corresponding continuous-time dynamical system converges to the Nash equilibrium associated with that incentive mechanism.
Assumption 3.2.
For any incentive mechanism $p,\tilde{p}$, consider the following continuous time dynamical systems associated with ($x$-update) and ($\tilde{x}$-update), respectively,
@@
The equilibrium $x^{\ast}(p)$ (resp. $\tilde{x}^{\ast}(\tilde{p})$) is the global asymptotically stable fixed point of the dynamics ($x$-dynamics) (resp. ($\tilde{x}$-dynamics)).
Using the standard approximation arguments from [Borkar, 2009] it is evident that Assumption 3.1 and 3.2 are sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms.
Against the preceding backdrop, in the following result, we characterize the convergence conditions for the coupled strategy and incentive updates.
@@
Consider an atomic game $G$. Suppose that Assumptions 2.1, 3.1, 3.2 and at least one of the following conditions holds:
  • (C1)
    $\frac{\partial e_{i}(x^{\ast}(p))}{\partial p_{j}}>0$ for all $p\in\mathbb{R}^{n}$ and all $i\neq j$. Additionally, at least one of the following conditions hold:
    • (i)
      If $e_{i}(x^{\ast}(0))\geqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{+}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\leqslant 0$ for every $i\in\mathcal{I}$. Moreover, $x_{1}\in X,p_{1}\in\mathbb{R}^{|\mathcal{I}|}_{+}$.
    • (ii)
      If $e_{i}(x^{\ast}(0))\leqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{-}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\geqslant 0$ for every $i\in\mathcal{I}$. Moreover, $x_{1}\in X,p_{1}\in\mathbb{R}^{|\mathcal{I}|}_{-}$.
  • (C2)
    There exists a set $\textsf{dom}(V)\subset\mathbb{R}^{|\mathcal{I}|}$ and a continuously differentiable function $V:\textsf{dom}(V)\rightarrow\mathbb{R}_{+}$ such that $V(p^{\dagger})=0$ and $V(p)>0$ for all $p\neq p^{\dagger}$. Moreover:
    $\nabla V(p)^{\top}\left(e(x^{\ast}(p))-p\right)<0\quad\forall\ p\neq p^{\dagger}.$
Additionally, let $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$. Then there exist positive scalars $\bar{r},\bar{\alpha},\bar{\beta},\bar{\gamma}$ such that the discrete-time updates ($x$-update)-($p$-update) in $G$ when initialized at $p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$, $x_{0}\in\mathcal{B}_{\bar{r}}(x^{\ast}(p_{0}))$, with step-sizes such that
$\sup_{k\in\mathbb{N}}\beta_{k}/\gamma_{k}\leqslant\bar{\alpha},\quad\sup_{k\in\mathbb{N}}\beta_{k}\leqslant\bar{\beta},\quad\sup_{k\in\mathbb{N}}\gamma_{k}\leqslant\bar{\gamma},$ (18)
satisfy [^2] The results in this paper hold even if the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) are perturbed with martingale difference noise [Borkar and Pattathil, 2018].
$\lim_{k\to\infty}(x_{k},p_{k})=({x}^{\dagger},p^{\dagger}).$ (19)
The same argument holds for non-atomic game $\tilde{G}$ under Assumptions 2.2, 3.1, and 3.2.
Proof.
Under Assumption 3.1, the two-timescale approximation theory [Borkar, 1997, Borkar and Pattathil, 2018] suggests that the strategy update ($x$-update) is a fast transient while the incentive update ($p$-update) is a slow component. This separation in timescales allow us to study the convergence of these updates in two stages. First, we study the convergence of fast strategy updates, for every fixed value of incentive. Second, we study the convergence of slow incentive updates, assuming that the fast strategy updates have converged to the equilibrium.
@@
$x_{k+1}$ $=x_{k}+\gamma_{k}\left(f(x_{k},p_{k})-x_{k}\right)$ (20)
@@
Since $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$ and $\lim_{k\rightarrow\infty}\beta_{k}/\gamma_{k}=0$ (cf. Assumption 3.1), the term $\frac{\beta_{k}}{\gamma_{k}}\left(e(x_{k})-p_{k}\right)$ in (20) goes to zero as $k\rightarrow\infty$. Consequently, leveraging the standard approximation arguments [Borkar, 2009, Lemma 1, Section 2.2], we conclude that the asymptotic behavior of the updates in (20) is same as that of the following dynamical system
@@
Next, to study the convergence of slow incentive updates, we re-write ($p$-update) as follows
@@
We will show that $(p_{k})_{k\in\mathbb{N}}$ will asymptotically follow the trajectories of the following continuous-time dynamical system:
@@
Note that $p^{\dagger}$ is the fixed point of the trajectories of dynamical system (23) (cf. Proposition 3.1). Conditions (C1) and (C2) in Proposition 3.3 are based on results from non-linear dynamical systems which ensure convergence of (23). In particular, (C1)-(i) (resp. (C1)-(ii)) builds on the cooperative dynamical systems theory [Hirsch, 1985], which ensures that $\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $\mathbb{R}_{-}^{|\mathcal{I}|}$) is positively invariant for (23) and $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $p^{\dagger}\in\mathbb{R}_{-}^{|\mathcal{I}|}$) is asymptotically stable. On the other hand, condition (C2) ensures the existence of a Lyapunov function that is strictly positive everywhere except at $p^{\dagger}$ and decreases along the flow of (23) ([Sastry, 2013]). Either condition guarantees the convergence of the trajectories of (23) to $p^{\dagger}$.
Let $D^{\dagger}$ denote the domain of attraction of $p^{\dagger}$ for the dynamical system (23). From converse Lyapunov theorem [Scheidegger, 1964], we know that there exists a continuously differentiable function $\bar{V}:D^{\dagger}\rightarrow\mathbb{R}_{+}$ such that $\bar{V}(p^{\dagger})=0$, $\bar{V}(p)>0$ for all $p\in D^{\dagger}\backslash\{p^{\dagger}\}$ and $\bar{V}(p)\rightarrow\infty$ as $p\rightarrow\textsf{boundary}(D^{\dagger})$. For any $r>0$, define $\bar{V}_{r}=\{p\in\textsf{dom}(\bar{V}):\bar{V}(p)\leqslant r\}$ to be a sub-level set of $\bar{V}$. There exists $0<\bar{r}^{\prime}<\bar{r}$ such that $\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})\subsetneq\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$. Additionally, define $\ell_{0}=0,\ell_{k}=\sum_{i=1}^{k}\beta_{i}$ and $L_{k}=\ell_{n(k)}$ where $n(0)=0$, and
@@
Consequently, using triangle inequality, it holds that
$\|p_{k}-p^{\dagger}\|\leqslant\kappa+\|\bar{\textbf{p}}^{(0)}(\ell_{k})-p^{\dagger}\|.$ (25)
Since $\bar{V}$ is a Lyapunov function of (23) and $\bar{\textbf{p}}^{(0)}(0)=p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$, there exists $\bar{k}\in\mathbb{N}$ such that for all $k\geqslant\bar{k}$, $\bar{\textbf{p}}^{(0)}(\ell_{k})\in\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})$. If we choose $\kappa<\bar{r}-\bar{r}^{\prime}$ then, from (25), it holds that for all $k\geqslant\bar{k}$, $p_{k}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$. Therefore, if we choose $T\geqslant\bar{k}$ in (24), it holds that
@@
Define $\hat{p}:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that, for every $k\in\mathbb{N}$, $\hat{p}(\ell_{k})=p_{k}$ with linear interpolation on $[\ell_{k},\ell_{k+1}]$. Using the standard approximation arguments from [Borkar, 2009, Chapter 6], it holds that333 For any $T\geqslant\bar{k}$ and $\delta>0$, there exists $k(\delta)$ such that $\hat{p}(\ell_{k(\delta)}+\cdot)$ form a “$(T,\delta)$” perturbation (cf. [Borkar, 1997]) of (23).
@@
Using (21) and Assumption 3.1, we conclude that RHS in the above equation goes to zero as $k\rightarrow\infty$. Finally, using (26), (3.2) and [Borkar, 1997, Lemma 2.1] we conclude that $p_{k}\rightarrow p^{\dagger}$ as $k\rightarrow\infty$. ∎

### v3

@@ 3.1 Fixed point analysis
$\langle\nabla\tilde{\Phi}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger})),\tilde{x}-{\tilde{x}^{\ast}}(\tilde{p}^{\dagger})\rangle\geqslant 0,\quad\forall\ \tilde{x}\in\tilde{X}.$ (17)
Comparing (17) with the first order necessary and sufficient conditions of optimality of social cost function, we note that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is the minimizer of the social cost function $\tilde{\Phi}$. This implies that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})={\tilde{x}}^{\dagger}$, since ${\tilde{x}}^{\dagger}$ is the unique minimizer of the social cost function $\tilde{\Phi}$ under Assumption 2.2.
Finally, we show that the set $P^{\dagger}$ is singleton. We prove this via contradiction. Suppose that $P^{\dagger}$ contains two element $p^{\dagger}_{1},p^{\dagger}_{2}$, and both align the Nash equilibrium with social optimality. Then, $x^{\dagger}=x^{\ast}(p^{\dagger}_{1})=x^{\ast}(p^{\dagger}_{2})$. From (LABEL:eq:_PDagger), we know that $p^{\dagger}_{1}=e(x^{\ast}(p^{\dagger}_{1}))$ and $p^{\dagger}_{2}=e(x^{\ast}(p^{\dagger}_{2}))$. Thus, we must have $p^{\dagger}_{1}=e(x^{\dagger})=p^{\dagger}_{2}$, which implies that $P^{\dagger}$ is a singleton. The proof of uniqueness of $\tilde{P}^{\dagger}$ follows analogously. $\square$
@@ 3.2 Convergence to optimal incentive mechanism
In this subsection, we provide sufficient conditions for the convergence of strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update). Before presenting the convergence result, we first introduce two assumptions.
@@
The step sizes in ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) satisfy the following conditions:
@@
    $\lim_{k\to\infty}{\beta_{k}}/{\gamma_{k}}=0$.
Assumption 3.1-(i) is a standard assumption on step sizes that allows us to analyze the convergence properties of the discrete-time learning updates through that of a continuous-time dynamical system [Borkar, 2009]. Assumption 3.1-(ii) ensures that the incentive update evolves on a slower timescale than the players’ strategy updates [Borkar, 2009, Lakshminarayanan and Bhatnagar, 2017]. Any step sizes of the form $\gamma_{k}=k^{-a}$ and $\beta_{k}=k^{-b}$ with $0.5<a<b\leqslant 1,$ satisfy Assumption 3.1.
Assumption 3.1 has been adopted in several previous works on adaptive incentive design (e.g., [Li et al., 2024, Chandak et al., 2024]). Under Assumption 3.1, the strategy update ($x$-update) represents a fast transient, whereas the incentive update ($p$-update) is a slow component. To analyze such discrete-time updates, we employ techniques from two-timescale approximation theory [Borkar, 2009, Borkar and Pattathil, 2018, Chandak et al., 2024], which allows us to analyze the convergence of the strategy and incentive updates separately. An intermediate step in this process is to ensure that, for every $p,\tilde{p}$, the trajectories of the following continuous-time strategy dynamics globally converge (cf. [Borkar, 2009, Borkar and Pattathil, 2018, Chandak et al., 2024]):
@@
In this work, we do not focus on analyzing the convergence of ($x$-dynamics)-($\tilde{x}$-dynamics). Instead, we assume any off-the-shelf convergent strategy update that satisfies the following assumption:
Assumption 3.2.
For any incentive mechanism $p$ (resp. $\tilde{p}$), the Nash equilibrium $x^{\ast}(p)$ (resp. $\tilde{x}^{\ast}(\tilde{p})$) is the globally asymptotically stable fixed point of the continuous-time dynamical system ($x$-dynamics) (resp. ($\tilde{x}$-dynamics)).
Assumption 3.2 is satisfied for a variety of strategy updates in various games. This includes the best-response and fictitious play strategy update in zero-sum and potential games [Hofbauer and Sorin, 2006, Benaïm et al., 2005, Swenson et al., 2018, Leslie and Collins, 2006], and gradient-based strategy update in continuous games [Mazumdar et al., 2020, Mertikopoulos et al., 2024].
Our goal here is to characterize conditions under which the coupled strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) converge. Before stating the convergence results, we define two notions of convergence.
Definition 3.2.
We say that the coupled strategy and incentive updates ($x$-update)-($p$-update)
  • (i)
    globally converges to the fixed point $(x^{\dagger},p^{\dagger})$ if, for any initial condition $p_{0}\in\mathbb{R}^{|\mathcal{I}|}$ and $x_{0}\in X$, and any selection of step sizes that satisfy Assumption 3.1, the discrete-time updates ($x$-update)-($p$-update) asymptotically converge to $(x^{\dagger},p^{\dagger})$.
  • (ii)
    locally converges to the fixed point $(x^{\dagger},p^{\dagger})$ if there exist positive scalars $\bar{r},\bar{\alpha},\bar{\beta},\bar{\gamma}$ such that, when $p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$ and $x_{0}\in\mathcal{B}_{\bar{r}}(x^{\ast}(p_{0}))$, the step sizes satisfy Assumption 3.1 and the following condition:
    $\sup_{k\in\mathbb{N}}\frac{\beta_{k}}{\gamma_{k}}\leqslant\bar{\alpha},\quad\sup_{k\in\mathbb{N}}\beta_{k}\leqslant\bar{\beta},\quad\sup_{k\in\mathbb{N}}\gamma_{k}\leqslant\bar{\gamma},$ (18)
    then the discrete-time updates ($x$-update)–($p$-update) asymptotically converge to $({x}^{\dagger},p^{\dagger})$.
Local and global convergence are analogously defined for the updates ($\tilde{x}$-update)-($\tilde{p}$-update) in non-atomic games.
@@
Consider the atomic game $G$ with discrete-time update ($x$-update)-($p$-update) that satisfy Assumptions 3.1 and 3.2. The following requirements provide sufficient conditions for ($x$-update)-($p$-update) to locally converge to the fixed point [^4] This result holds even if the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) are perturbed with square-integrable martingale difference noise [Borkar and Pattathil, 2018]. $(x^{\dagger},p^{\dagger})$ in the sense of Definition 3.2:
  1. (R1)
    $p^{\dagger}$ is a locally asymptotically stable equilibrium of the following continuous-time dynamical system:
    $\dot{\textbf{p}}(t)=e(x^{\ast}(\textbf{p}(t)))-\textbf{p}(t).$ (19)
  2. (R2)
    The trajectories of the discrete-time updates satisfy the boundedness condition:
    $\sup_{k\in\mathbb{N}}\big{(}\|x_{k}\|+\|p_{k}\|\big{)}<+\infty.$
Furthermore, the sufficient conditions for ($x$-update)-($\tilde{p}$-update) to globally converge to the fixed point $(x^{\dagger},p^{\dagger})$ in the sense of Definition 3.2 are (R1’) and (R2), where
  1. (R1’)
    $p^{\dagger}$ is a globally asymptotically stable equilibrium of the continuous-time dynamical system:
    $\dot{\textbf{p}}(t)=e(x^{\ast}(\textbf{p}(t)))-\textbf{p}(t).$ (20)
Analogous result holds for the non-atomic game $\tilde{G}$.
Proposition 3.3 states two sets of generic conditions that can be verified when studying convergence in any specific game. In particular, by leveraging results from nonlinear dynamical systems theory, (R1) (or (R1’)) can be verified by showing the existence of a Lyapunov function [Sastry, 2013] or by establishing that the dynamical system is cooperative [Hirsch, 1985]; see Lemma D.1 in Appendix D. Additionally, (R2) holds in any game with a compact strategy set. In games with an unbounded strategy set, (R2) can be verified by analyzing the global convergence of continuous-time (‘scaled’) strategy and incentive dynamics [Lakshminarayanan and Bhatnagar, 2017, Theorem 10]. In Section 4, we verify that the conditions in Proposition 3.3 are satisfied in atomic aggregative games and non-atomic routing games. Proof of Proposition 3.3. Assumption 3.1-(ii), allow us to study the convergence of ($x$-update)-($p$-update) in two stages [Borkar, 1997, Borkar and Pattathil, 2018]. First, we study the convergence of fast strategy updates, for every fixed value of incentive. Second, we study the convergence of slow incentive updates, assuming that the fast strategy updates have converged to the equilibrium.
@@
$x_{k+1}$ $=x_{k}+\gamma_{k}\left(f(x_{k},p_{k})-x_{k}\right),$ (21)
@@
Since $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$ (cf. requirement (R2)) and $\lim_{k\rightarrow\infty}\beta_{k}/\gamma_{k}=0$ (cf. Assumption 3.1), the term $\frac{\beta_{k}}{\gamma_{k}}\left(e(x_{k})-p_{k}\right)$ in (21) goes to zero as $k\rightarrow\infty$. Consequently, leveraging the standard approximation arguments [Borkar, 2009, Lemma 1, Section 2.2], we conclude that the asymptotic behavior of the updates in (21) is same as that of the following dynamical system
@@
Next, to study the convergence of the slow incentive updates, we re-write ($p$-update) as follows
@@
We will show that $(p_{k})_{k\in\mathbb{N}}$ will asymptotically follow the trajectories of the following continuous-time dynamics:
@@
Note that $p^{\dagger}$ is the fixed point of the trajectories of the dynamical system (19) (cf. Proposition 3.1). Requirement (R1) in Proposition 3.3 ensures convergence of (19).
Let $D^{\dagger}$ denote the domain of attraction of $p^{\dagger}$ for the dynamical system (19). From the converse Lyapunov theorem [Scheidegger, 1964], we know that there exists a continuously differentiable function $\bar{V}:D^{\dagger}\rightarrow\mathbb{R}_{+}$ such that $\bar{V}(p^{\dagger})=0$, $\bar{V}(p)>0$ for all $p\in D^{\dagger}\backslash\{p^{\dagger}\}$ and $\bar{V}(p)\rightarrow\infty$ as $p\rightarrow\textsf{boundary}(D^{\dagger})$. For any $r>0$, define $\bar{V}_{r}=\{p\in\textsf{dom}(\bar{V}):\bar{V}(p)\leqslant r\}$ to be a sub-level set of $\bar{V}$. There exists $0<\bar{r}^{\prime}<\bar{r}$ such that $\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})\subsetneq\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$. Additionally, define $t_{0}=0,t_{k}=\sum_{i=1}^{k}\beta_{i}$ and $L_{k}=t_{n(k)}$ where $n(0)=0$, and
@@
Consequently, using the triangle inequality, it holds that
$\|p_{k}-p^{\dagger}\|\leqslant\kappa+\|\bar{\textbf{p}}^{(0)}(t_{k})-p^{\dagger}\|.$ (26)
Since $\bar{V}$ is a Lyapunov function of (19) and $\bar{\textbf{p}}^{(0)}(0)=p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$, there exists $\bar{k}\in\mathbb{N}$ such that for all $k\geqslant\bar{k}$, $\bar{\textbf{p}}^{(0)}(t_{k})\in\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})$. If we choose $\kappa<\bar{r}-\bar{r}^{\prime}$ then, from (26), it holds that for all $k\geqslant\bar{k}$, $p_{k}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$. Therefore, if we choose $T\geqslant\bar{k}$ in (25), it holds that
@@
Define $\hat{p}:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that, for every $k\in\mathbb{N}$, $\hat{p}(t_{k})=p_{k}$ with linear interpolation on $[t_{k},t_{k+1}]$. Using the standard approximation arguments from [Borkar, 2009, Chapter 6], it holds that555 For any $T\geqslant\bar{k}$ and $\delta>0$, there exists $k(\delta)$ such that $\hat{p}(t_{k(\delta)}+\cdot)$ form a “$(T,\delta)$” perturbation (cf. [Borkar, 1997]) of (19).
@@
Using (22) and Assumption 3.1, we conclude that RHS in the above equation goes to zero as $k\rightarrow\infty$. Finally, using (27), (3.2) and [Borkar, 1997, Lemma 2.1], we conclude that $p_{k}\rightarrow p^{\dagger}$ as $k\rightarrow\infty$. $\square$